Find the equation of the tangent to the parabola y^2=12x at the point (2, 5)
Answer by Alan3354(69443)

Find the equation of the tangent to the parabola y^2=12x at the point (2, 5)
-------------
y^2=12x
2ydy = 12dx
dy/dx = 6/y
y'(5) = 6/5
----------------
y-5 = (6/5)*(x-2) is the tangent at (2,5)